# The inner structure and thermodynamics of a thin accretion disc

### v1

@@ Abstract
Using three-dimensional general relativistic magnetohydrodynamic simulations with electron and proton thermodynamics, we probe the inner radial and vertical structure of weakly magnetized geometrically thin accretion discs around rapidly spinning black holes. We find that the thin, cold disc transitions to a thick, hot accretion flow at a radius dependent on the mass accretion rate. At high accretion rates, the disc truncates close to the innermost stable circular orbit $r\approx 2r_{g}$, demonstrating that even in the canonical thin disc model, the plunging region should be treated with two-temperature physics. At intermediate accretion rates, the transition radius moves outward by a factor of two to $r\approx 5r_{g}$, forming a radiatively inefficient inner flow. The simulations also reveal extended cooling along the surface of the disc out to $\sim 10r_{g}$, with 40% of the total cooling at intermediate accretion rates occurring above the disc body. These results have implications for X-ray binary state transitions and the physical origin of the X-ray corona.
@@
pubyear: 2025pagerange: The inner structure and thermodynamics of a thin accretion discB
@@ 1 Introduction
Geometrically thin, optically thick accretion discs (Shakura & Sunyaev, 1973; Novikov & Thorne, 1973) are generally thought to power luminous black hole X-ray binaries (XRBs) and active galactic nuclei (AGN). Thin disc models assume that the disc electrons efficiently radiate the viscously dissipated energy that heats both electrons and protons. The protons in thin disc models cool due to efficient Coulomb collisions with electrons enabled by high densities. Standard models often assume that the flow truncates abruptly at the innermost stable circular orbit (ISCO), with no dissipation or radiation inside of that radius. However, observations (Fabian et al., 2020), analytic models (Gammie, 1999; Agol & Krolik, 2000), and numerical simulations (Noble et al., 2009; Zhu et al., 2012; Avara et al., 2016; Dhang et al., 2025) have suggested that magnetic torques may result in dissipation and non-negligible amounts of radiation coming from within the plunging region. These models have implicitly assumed that the electrons and protons remain well-coupled within the ISCO, despite hints to the contrary (Zhu et al., 2012). Analytic estimates showed that electrons and protons may decouple within the ISCO, potentially explaining the $>10$ keV tail observed in the soft state of XRBs (Hankla et al., 2022).
@@
Numerical simulations have attempted to realize global models of a thin accretion disc using general relativistic ideal magnetohydrodynamics (GRMHD). Since the disc cools by radiation, such simulations have either used approximate forms of radiative transfer (Liska et al., 2022) or idealized the cooling as a source term that removes thermal energy from the gas (Shafee et al., 2008; Noble et al., 2009; Penna et al., 2010). These prescriptions often assume perfect coupling between electrons and ions by treating the gas as a single fluid. In some cases, radiation is then postprocessed onto the disc to calculate predictions for XRB spectra (Kinch et al., 2020). Two-temperature thin discs that evolve electrons as a separate fluid (Ressler et al., 2015) have been mainly studied in the context of strong magnetic fields at high accretion rates, where magnetic torques cause the thin disc to transition to a hot flow at small radii (Liska et al., 2022).
@@
Proton-electron coupling through Coulomb collisions plays an important role in accretion-rate-dependent models for XRB state transitions. In the “disc truncation” model, the thin accretion disc transitions to a hot accretion flow at radii $\gg r_{\rm ISCO}$ at intermediate accretion rates (Esin et al., 1997; Done et al., 2007). Whereas the soft state of XRBs is well described by a cold, thin disc, the hard state likely originates from photons Compton scattering off a geometrically thick, optically-thin accretion flow (Narayan & Yi, 1994; Yuan & Narayan, 2014). Geometrically thick, optically-thin accretion flows have low densities, inefficient Coulomb coupling, and therefore very little proton cooling. In these discs, the electron cooling does not influence the disc structure, motivating the evolution of a single-temperature flow whose proton-to-electron temperature ratio and thus radiation is calculated during postprocessing (Mościbrodzka et al., 2009). GRMHD simulations that treat radiation through Monte Carlo methods have previously suggested that electron dynamics become important at accretion rates higher than expected by analytic models (Ryan et al., 2017), potentially even collapsing the hot flow to a thin disc (Dexter et al., 2021).
@@
Because of the difficulty in treating both radiation and thermodynamics, the transition from thick to thin disc has only been studied in local context or with a pre-determined truncation radius. Radiation in particular becomes difficult to model in these intermediate accretion rate regimes. Monte Carlo methods become computationally intractable as the photon optical depth increases the number of scatterings, whereas treating radiation as a fluid in optically-thin regions may not accurately treat anisotropic and non-local Compton cooling. Therefore, numerical simulations have attempted to access this radial transition by putting in the transition by hand and examining the resulting structure (Hogg & Reynolds, 2017, 2018). A complementary approach is to examine the vertical structure alone using local shearing box simulations. Although these simulations have found vertical decoupling of protons and electrons (Bambic et al., 2024), by their nature they cannot probe a radial transition in the accretion disc structure.
@@ 2 Methods
The simulations in this work use nubhlight (Miller et al., 2019), an extension of the public GRMHD code ebhlight (Ryan et al., 2015, 2017), which solves the general relativistic ideal magnetohydrodynamic equations using the HARM scheme (Gammie et al., 2003; Noble et al., 2006). An electron fluid, evolved separately from the background fluid (Ressler et al., 2015), heats via Coulomb collisions and grid-scale dissipation. The viscously dissipated energy partitions into electron and proton heating according to particle-in-cell simulations of 2D magnetic reconnection with a moderate guide field (Werner et al., 2018). This prescription gives a fraction $\delta_{e}$ between 1/4 and 1/2 of the dissipated viscous energy to electrons according to the magnetization of the plasma in a given simulation cell. The cooling function described below affects only this electron fluid, in contrast to single-fluid GRMHD cooling functions.
@@ 2.3 GRMHD Simulations
The simulations have $[320,256,160]$ cells in the $[r,\theta,\phi]$ directions. Unless otherwise noted, the simulation data are time-averaged over the last $2000r_{g}/c$, with the final times given in Table 1.
@@ 2.4 Diagnostics
To exclude the floor-dominated region within the jet, the integration is performed over regions with plasma magnetization $\sigma\equiv b_{\mu}b^{\mu}/\rho<1$.
The mass accretion rate is defined as the mass flux at the event horizon:
$\dot{M}\equiv\dot{M}(r_{\rm EH})=-\iint\rho u^{r}\sqrt{-g}~{}{\rm d}\theta~{}{\rm d}\phi,$ (6)
and we refer to its value time-averaged over the last $2\times 10^{3}r_{g}/c$, as outlined in Table 1.
@@
$\tau_{es}(r)=\frac{1}{2}\int_{0}^{2\pi}~{}{\rm d}\phi\left(\int_{0}^{\pi/2}\kappa\rho(r,\theta,\phi)~{}\sqrt{g_{\theta\theta}}{\rm d}\theta\right.$ (8)
$\left.+\int_{\pi}^{\pi/2}\kappa\rho(r,\theta,\phi)~{}\sqrt{g_{\theta\theta}}{\rm d}\theta\right)$ (9)
@@ 3.1 The Plunging Region of a Thin Disc
We begin with a case study comparing a low accretion rate simulation where Coulomb collisions are negligible (M6) to a high accretion rate simulation (M0) where Coulomb collisions collapse the initially thick torus into a thin disc with the target scale height.
@@
In contrast, the thin accretion disc’s structure changes as Coulomb collisions become important in most regions of the disc (Fig. 1 right column). For large radii, the proton cooling timescale due to Coulomb collisions equals the proton heating timescale due to viscous dissipation, and both timescales happen efficiently before material falls onto the black hole: $t^{i}_{\rm visc}\simeq t^{i}_{\rm Coul}\ll t_{\rm infall}$. Similarly, the electron heating and cooler timescales at large radii are also faster than the infall time.
@@ 3.2 Dependence of disc Properties on Accretion Rate
We next investigate the impact of Coulomb collisions on the structure of the disc by holding the electron target temperature at a constant value of $10^{10}~{}{\rm K}$ and varying the scaling of the mass unit, thereby adjusting the accretion rate. Because the Coulomb energy exchange rate depends quadratically on the mass unit, adjusting the accretion rate will increase the importance of Coulomb collisions throughout the disc. In this section, we explore where Coulomb collisions become important enough to affect the disc structure in the radial and vertical directions. All accretion rates are relatively constant over time; the time-averaged values are listed in Table 1.
@@ 3.2.3 Timescales
To understand the origin of the thermodynamics and structure presented in Figs. 4 and 5, Fig. 6 plots the relevant timescales in the system (Eqns. 11 and 12). These timescales are normalized to the infall time (Eq. 10) such that lines above 1 show physical processes that cannot operate before the gas accretes. The timescales are separated into proton (top) and electron (bottom) to demonstrate how the electron cooling function affects the gas structure by influencing protons through Coulomb collisions.
@@ 4.1 Disc Truncation Radius Depends on Accretion Rate
Although the target temperature in these simulations is larger than expected from Compton scattering, the presence of the decoupled flow over a factor of 100 in accretion rate suggests that the truncation is a robust result. The exact accretion rates listed in these simulations may not remain the same with a lower target temperature since the proton temperature and thus density would also be affected by a change in target temperature. However, the presence of the truncation over such a large range in accretion rate suggests that models with more accurate temperatures would also result in truncation.
@@ 4.1.1 Implications for XRB State Transitions
Notably, the disc truncation presented in this simulation suite occurs due to Coulomb collisions where gas pressure dominates the disc structure. In accretion discs with stronger magnetic fields, the magnetic fields likely play a significant role in truncating the disc (Liska et al., 2022). XRB state transitions are likely governed by a combination of accretion rate changes and magnetic field advection (Begelman & Armitage, 2014).
@@ 4.1.2 Implications For Particle Acceleration within the ISCO
The truncation of a thin disc into a two-temperature flow at high accretion rates has implications for nonthermal particle acceleration within the ISCO. Because of the drop in density within the ISCO by two orders of magnitude, the collision times between electrons also drop. These conditions could be sufficient to produce the hard X-ray tail $>10~{}{\rm keV}$ observed in the soft state of XRBs (McConnell et al., 2002); see Hankla et al. (2022).
@@ 4.3.1 Vertical Extent of the Two-Temperature Regions
Fig. 11 shows the distribution of luminosity from the electron cooling function. Notably, a significant amount of cooling originates from above the disc out to $\sim 10r_{g}$ in a sandwich-esque configuration. Within the truncated region, the emission clusters close to the equatorial plane rather than becoming perfectly spherical. We again note that the emission outside the contour for $T_{p}/T_{e}=10$ is likely highly nonthermal. The amount of radiation from above the disc body is similar to the amount found in single-temperature simulations of thin, strongly magnetized accretion discs (Scepi et al., 2024) and local, two-temperature simulations (Bambic et al., 2024). These findings have implications for the shape of the X-ray corona as probed by spectral-timing studies and X-ray polarimetry.
4.4 Other Considerations

### v2

# The inner structure and thermodynamics of thin accretion discs

@@ Abstract
Using three-dimensional general relativistic magnetohydrodynamic simulations with electron and proton thermodynamics and an electron cooling function, we probe the inner radial and vertical structure of weakly magnetized geometrically thin accretion discs around rapidly spinning black holes. We find that the thin, cold disc transitions to a thick, hot accretion flow at a radius dependent on the mass accretion rate due to proton-electron Coulomb decoupling. At high accretion rates, the disc truncates close to the innermost stable circular orbit $r\approx 2r_{g}$, demonstrating that even in the canonical thin disc model, the plunging region should be treated with two-temperature physics. At intermediate accretion rates, the transition radius moves outward by a factor of two to $r\approx 5r_{g}$, forming a radiatively inefficient inner flow. The simulations also reveal extended cooling along the surface of the disc out to $\sim 10r_{g}$, with 40% of the total cooling at intermediate accretion rates occurring above the disc body. Two-temperature effects also impact the emission from within the plunging region of the black hole, leading to less thermal emission than predicted by single-temperature models. These results have implications for X-ray binary state transitions, the physical origin of the X-ray corona, and spin measurements that rely on determining the location of the innermost stable circular orbit.
@@
pubyear: 2025pagerange: The inner structure and thermodynamics of thin accretion discsB
@@ 1 Introduction
Geometrically thin, optically thick accretion discs (Shakura & Sunyaev, 1973; Novikov & Thorne, 1973) are generally thought to power luminous black hole X-ray binaries (XRBs) and active galactic nuclei (AGN). Thin disc models assume that the disc electrons efficiently radiate the viscously dissipated energy that heats both electrons and protons. The protons in thin disc models cool due to efficient Coulomb collisions with electrons enabled by high densities. Standard models often assume that the flow truncates abruptly at the innermost stable circular orbit (ISCO), with no dissipation or radiation inside of that radius. However, observations (Fabian et al., 2020), analytic models (Gammie, 1999; Agol & Krolik, 2000), and numerical simulations (Noble et al., 2009; Zhu et al., 2012; Avara et al., 2016; Scepi et al., 2024a; Dhang et al., 2025) have suggested that magnetic torques may result in dissipation and non-negligible amounts of radiation coming from within the plunging region. These models have implicitly assumed that the electrons and protons remain well-coupled within the ISCO, despite hints to the contrary (Zhu et al., 2012). Analytic estimates showed that electrons and protons may decouple within the ISCO, potentially explaining the $>10$ keV tail observed in the soft state of XRBs (Hankla et al., 2022).
@@
Numerical simulations have attempted to realize global models of a thin accretion disc using general relativistic ideal magnetohydrodynamics (GRMHD). Since the disc cools by radiation, such simulations have either used approximate forms of radiative transfer (Liska et al., 2022) or idealized the cooling as a source term that removes thermal energy from the gas (Shafee et al., 2008; Noble et al., 2009; Penna et al., 2010). These prescriptions often assume perfect coupling between electrons and ions by treating the gas as a single fluid. In some cases, radiation is then postprocessed onto the disc to calculate predictions for XRB spectra (Kinch et al., 2020). Two-temperature thin discs that evolve electrons as a separate fluid (Ressler et al., 2015) have been mainly studied in the context of strong magnetic fields at high accretion rates, where magnetic torques cause the thin disc to transition to a hot flow at small radii (Liska et al., 2022). These magnetically-arrested (MAD) discs can appear observationally truncated at a radius not necessarily related to the radius at which the disc becomes MAD (Scepi et al., 2024b).
@@
Proton-electron coupling through Coulomb collisions plays an important role in accretion-rate-dependent models for XRB state transitions. In the “disc truncation” model, the thin accretion disc transitions to a hot accretion flow at radii $\gg r_{\rm ISCO}$ at intermediate accretion rates (Esin et al., 1997; Done et al., 2007). Whereas the soft state of XRBs is well described by a cold, thin disc, the hard state likely originates from photons Compton scattering off a geometrically thick, optically-thin accretion flow (Narayan & Yi, 1994; Yuan & Narayan, 2014). Geometrically thick, optically-thin accretion flows have low densities, inefficient Coulomb coupling, and therefore very little proton cooling. In these discs, the electron cooling does not influence the disc structure, motivating the evolution of a single-temperature flow whose proton-to-electron temperature ratio and thus radiation is calculated during postprocessing (Mościbrodzka et al., 2009). GRMHD simulations that treat radiation through Monte Carlo methods have previously suggested that electron dynamics become important at accretion rates higher than expected by analytic models (Ryan et al., 2017), potentially collapsing the hot flow to a thin disc (Dexter et al., 2021).
@@
Because of the difficulty in treating both radiation and thermodynamics, the transition from thick to thin disc has only been studied in local context or with a pre-determined truncation radius. Radiation becomes difficult to model in these intermediate accretion rate regimes. Monte Carlo methods become computationally intractable as the photon optical depth increases the number of scatterings, whereas treating radiation as a fluid in optically-thin regions may not accurately treat anisotropic and non-local Compton cooling. Therefore, numerical simulations have attempted to access this radial transition by putting in the transition by hand and examining the resulting structure (Hogg & Reynolds, 2017, 2018). A complementary approach is to examine the vertical structure alone using local shearing box simulations. Although these simulations found vertical decoupling of protons and electrons (Bambic et al., 2024), by their nature they cannot probe a radial transition in the accretion disc structure.
@@ 2 Methods
The simulations in this work use nubhlight (Miller et al., 2019), an extension of the public GRMHD code ebhlight (Ryan et al., 2015, 2017), which solves the general relativistic ideal magnetohydrodynamic equations using the HARM scheme (Gammie et al., 2003; Noble et al., 2006). An electron fluid, evolved separately from the background fluid (Ressler et al., 2015), heats via Coulomb collisions and grid-scale dissipation. The viscously dissipated energy partitions into electron and proton heating according to particle-in-cell simulations of 2D magnetic reconnection with a moderate guide field; see Werner et al. (2018) Eq. 3. This prescription gives a fraction $\delta_{e}$ between 1/4 and 1/2 of the dissipated viscous energy to electrons according to the magnetization of the plasma $\sigma_{i}$ in a given simulation cell: $\delta_{e}=0.25(1+\sqrt{\sigma_{i}/5/(2+\sigma_{i}/5)}$. The cooling function described below affects only this electron fluid, in contrast to single-fluid GRMHD cooling functions.
@@ 2.3 GRMHD Simulations
The simulations have $[320,256,160]$ cells in the $[r,\theta,\phi]$ directions. The simulation domain extends $2\pi$ radians in $\pi$, $\pi$ radians in $\theta$, and from within the event horizon out to $r=1000r_{g}$. Unless otherwise noted, the simulation data are time-averaged over the last $2000r_{g}/c$, with the final times given in Table 1. The simulations use the standard drift-frame floors: density and total internal energy density are added in the drift frame when their values dip beneath the floor values $\rho_{\rm floor}={\rm max}[b^{\mu}b_{\mu}/50,10^{-6}(r/r_{g})^{-2}]$ and $u_{\rm floor}={\rm max}[b_{\mu}b^{\mu}/2500,~{}10^{-8}(r/r_{g})^{-2\gamma}]$, respectively (Ressler et al., 2015, 2017). Throughout, we use code units unless otherwise specified.
@@ 2.4 Diagnostics
To exclude the floor-dominated region within the jet, the integration is performed over regions with magnetization $\sigma\equiv b_{\mu}b^{\mu}/\rho<1$.
The mass accretion rate is defined as the mass flux at the event horizon and we refer to its value time-averaged over the last $2\times 10^{3}r_{g}/c$, as outlined in Table 1:
$\dot{M}\equiv\dot{M}(r_{\rm EH})=-\iint\rho u^{r}\sqrt{-g}~{}{\rm d}\theta~{}{\rm d}\phi.$ (6)
@@
$\tau_{es}(r)=\frac{1}{2}\int_{0}^{2\pi}~{}{\rm d}\phi\left(\int_{0}^{\pi/2}\kappa_{\rm es}\rho(r,\theta,\phi)~{}\sqrt{g_{\theta\theta}}{\rm d}\theta\right.$ (8)
$\left.+\int_{\pi}^{\pi/2}\kappa_{\rm es}\rho(r,\theta,\phi)~{}\sqrt{g_{\theta\theta}}{\rm d}\theta\right)$ (9)
@@
Figure 1: Azimuthally-averaged density structure of the low accretion rate simulation M6 (left) and highest accretion rate simulation M0 (right). The white dotted contour shows $\tau_{\rm es}=10$, while the dark blue dash-dot and dashed lines show $T_{p}/T_{e}=2$ and 10, respectively. The red dashed line shows the equatorial ISCO radius. Time-averaged over the last $2000r_{g}/c$.
@@ 3.1 The Plunging Region of a Thin Disc
We begin with a case study comparing a low accretion rate simulation where Coulomb collisions are negligible (M6) to a high accretion rate simulation (M0) where Coulomb collisions collapse the initially thick torus into a thin disc with the target scale height. Fig. 1 shows the difference in density structure for these two simulations.
@@
In contrast, the thin accretion disc’s structure changes as Coulomb collisions become important in most regions of the disc (Fig. 2 right column). For large radii, the proton cooling timescale due to Coulomb collisions equals the proton heating timescale due to viscous dissipation, and both timescales happen efficiently before material falls onto the black hole: $t^{i}_{\rm visc}\simeq t^{i}_{\rm Coul}\ll t_{\rm infall}$. Similarly, the electron heating and cooling timescales at large radii are also faster than the infall time.
@@ 3.2 Dependence of Disc Properties on Accretion Rate
We next investigate the impact of Coulomb collisions on the structure of the disc by holding the electron target temperature at a constant value of $10^{10}~{}{\rm K}$ and varying the scaling of the mass unit, thereby adjusting the accretion rate. Because the Coulomb energy exchange rate depends quadratically on the mass unit, increasing the accretion rate will increase the importance of Coulomb collisions throughout the disc. In this section, we explore where Coulomb collisions become important enough to affect the disc structure in the radial and vertical directions. All accretion rates are relatively constant over time; the time-averaged values are listed in Table 1.
@@ 3.2.3 Timescales
To understand the origin of the thermodynamics and structure presented in Figs. 5 and 6, Fig. 7 plots the relevant timescales in the system (Eqns. 11 and 12). These timescales are normalized to the infall time (Eq. 10) such that lines above 1 show physical processes that cannot operate before the gas accretes. The timescales are separated into proton (top) and electron (bottom) to demonstrate how the electron cooling function affects the gas structure by influencing protons through Coulomb collisions.
@@ 4.1 Disc Truncation Radius Depends on Accretion Rate
Although the target temperature in these simulations is larger than expected from Compton scattering, the presence of the decoupled flow over a factor of 100 in accretion rate suggests that the truncation is a robust result. The exact accretion rates listed in these simulations may not remain the same with a lower target temperature since the proton temperature and thus density would also be affected by a change in target temperature. However, the presence of the truncation over such a large range in accretion rate suggests that models with more accurate temperatures would also result in truncation, at least within the ISCO. A simulation with a target temperature lowered by a factor of 100, yielding a density a factor of approximately 10 higher, will still decouple within the ISCO, as shown analytically (Hankla et al., 2022). Future work will examine more realistic target temperatures, including prescribing $H/r$ as constant (with $T_{e,\rm target}\propto T^{1/2}$) as is the case in radiation-pressure-dominated discs.
@@ 4.1.1 Implications for XRB State Transitions
Notably, the disc truncation presented in this simulation suite occurs due to Coulomb collisions where gas pressure dominates the disc structure, namely the disc does not become magnetically arrested (MAD). In accretion discs with stronger magnetic fields, the magnetic fields likely play a significant role in truncating the disc (Liska et al., 2022), although the radius that the disc observationally truncates could occur at some location within the MAD radius (Scepi et al., 2024b). XRB state transitions are likely governed by a combination of accretion rate changes and magnetic field advection (Begelman & Armitage, 2014).
@@ 4.1.2 Implications For Particle Acceleration within the ISCO
The truncation of a thin disc into a two-temperature flow at high accretion rates has implications for nonthermal particle acceleration within the ISCO. Because of the drop in density within the ISCO by two orders of magnitude, the collision times between electrons also drop, as seen in the Coulomb exchange time $t^{e,i}_{\rm Coul}$ in Figs. 2 and 7 and calculated analytically in Hankla et al. (2022). These conditions could be sufficient to produce the hard X-ray tail $>10~{}{\rm keV}$ observed in the soft state of XRBs (McConnell et al., 2002).
@@ 4.3.1 Vertical Extent of the Two-Temperature Regions
Fig. 12 shows the distribution of luminosity from the electron cooling function. Notably, a significant amount of cooling originates from above the disc out to $\sim 10r_{g}$ in a sandwich-esque configuration. Within the truncated region, the emission clusters close to the equatorial plane rather than becoming perfectly spherical. We again note that the emission outside the contour for $T_{p}/T_{e}=10$ is likely highly nonthermal. The amount of radiation from above the disc body is similar to the amount found in single-temperature simulations of thin, strongly magnetized accretion discs (Scepi et al., 2024a) and local, two-temperature simulations (Bambic et al., 2024). These findings have implications for the shape of the X-ray corona as probed by spectral-timing studies and X-ray polarimetry.
4.4 Impact of Two-temperature Physics on Black Hole Spin Measurements
A common method to measure black hole spin involves identifying the disc peak temperature and mapping that temperature to a specific radius related to the ISCO, which then provides the black hole spin; see Reynolds (2021). This thermal continuum fitting model is therefore extremely sensitive to the location of the maximum temperature. Two-temperature physics impacts the location of maximum temperature in two main ways.
First, at intermediate accretion rates, Coulomb decoupling leads to the transition away from a thin disc at radii larger than the ISCO (e.g. Fig. 11). For example, the M2 simulation’s ISCO would be observationally identified as having an ISCO around $4r_{g}$, rather than the true value of $\approx 2r_{g}$ for $a=0.9375$. Thus truncation due to two-temperature effects could lead to underestimating black hole spin for intermediate accretion rates.
Second, two-temperature effects reduce the impact of thermal emission from within the ISCO compared to single-temperature models. Single-temperature models have long demonstrated that thermal emission from within the ISCO can lead to misidentification of the ISCO location, thereby overestimating black hole spin (Penna et al., 2010). As we show in Sec. 3.3, including two-temperature effects reduces the efficiency of thermal emission from within the ISCO. Because Coulomb collisions become inefficient, protons heat the electrons less, leading to less electron cooling than in single-temperature models (Mummery et al., 2024, e.g.). At the highest accretion rate, the thermal emission at the ISCO is $\approx 75\%$ of the single-temperature prediction, dropping to $<50\%$ for lower accretion rates (Fig. 9). In addition, two-temperature physics shows that a significant amount of radiation from within the ISCO is optically thin (Fig. 10 central panel) rather than thermal. Emission from the plunging region therefore likely contains a significant nonthermal component rather than appearing purely thermal, as predicted analytically (Hankla et al., 2022) and suggested by detailed postprocessing of GRMHD simulations (Zhu et al., 2012).
4.5 Numerical Considerations